# Distributed Model Checking in Graphs Classes

### v1

@@ 5.3 Certifying Reductions on Graphs of Bounded Expansion
Definition 4.
Let $\varphi\in\text{FO}[\Lambda]$. For every labeled graph $(G,L)$ with $G\in\mathcal{G}$, let us assume that every node $u$ of $G$ is provided with the following values:
@@
    The set $\Lambda$, the formula $\varphi$, and the label $L(u)\subseteq\Lambda$;
@@
    A set $\widehat{\Lambda}$, a formula $\widehat{\varphi}$, and a label $\widehat{L}(u)\subseteq\widehat{\Lambda}$;
@@
    The encoding of a $p$-skeleton $\tau_{p}$ of $G$.
We say that $\varphi$ admits a distributedly certifiable reduction on $\mathcal{G}$ if there is a certification scheme using $\mathcal{O}(\log n)$-bits certificates for certifying that $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{L})$ is the reduction of $(\Lambda,\varphi,\tau_{p},L)$ on $G$.
In a way similar to what was done in Section 3, we show that every first-order formula over labeled graphs admits a distributedly certifiable reduction of $\mathcal{G}$. We fist establish this result for existential formulas.
Lemma 18.
Every existential formula in $\text{FO}[\Lambda]$ admits a distributed certifiable reduction on $\mathcal{G}$.
@@
Let $\varphi\in\text{FO}[\Lambda]$ be an existential formula. Consider that each node is provided with the values listed in 4. Let $\varphi(\overline{x})=\exists\overline{y}\,\psi(\overline{x},\overline{y})$, where $\psi$ is quantifier free. Every node checks that $p=|\overline{x}|+|\overline{y}|$. The scheme of Lemma 17 enables to certify that the encoding $\tau_{p}$ correctly defines a $p$-skeleton of $G$. Let $C_{p}$ be the $(p,f(p))$-treedepth coloring that defines $\tau_{p}$. The same construction as in Lemma 8 can then be used. More precisely, for evey $U\in\textsf{Col}(p)$, let $\widetilde{\Lambda}^{U}$, $\widetilde{\psi}^{U}$, $\widetilde{L}^{U}$, $\Lambda^{U}$, $\psi^{U}$, and $L^{U}$ be defined as in the proof of Lemma 8. Instead of computing these values, each node receives them from the prover as certificates. The values of $\widehat{\Lambda}$, $\widehat{\varphi}$, $\widetilde{\Lambda}^{U}$, $\widetilde{\psi}^{U}$, $\Lambda^{U}$, and $\psi^{U}$ do not depend on the input graph, and can be computed by each node. The labeling $\tilde{L}^{U}$ can be computed from $L$ by a 1-round algorithm in the BCLIQUE model. Hence, using Lemma 15, the correctness of each labeling $\tilde{L}^{U}$ can be certified with $\mathcal{O}(\log n)$-bit certificates. The correctness of each labeling $\widehat{L}^{U}$ can be certified using Lemma 16, using $\mathcal{O}(\log n)$-bit certificates. Finally, each node can check the correctness of $\widehat{L}$, and accept or reject accordingly. ∎
@@
Lemma 19.
Every formula in $\text{FO}[\Lambda]$ admits a distributed certifiable reduction on $\mathcal{G}$.
@@
Let $\varphi\in\text{FO}[\Lambda]$ be an existential formula. Let us assume that each node is provided with the values described in Definition 4, i.e. the tuples $(\Lambda,\varphi,L)$ and $(p,\hat{\Lambda},\hat{\varphi},\hat{L})$, and the encoding of a $p$-skeleton $\tau_{p}$. We follow the same induction as in Lemma 8, over the quantifier depth of $\varphi$. The base case is trivial as the construction requires no communication. Furthermore, if $\varphi=\neg\psi$ where $\psi$ admits a distributed certifiable reduction, then $\varphi$ admits a distributed certifiable reduction as well, as we described in Lemma 8.
It remains to consieder the case $\varphi(\overline{x})=\exists y\,\psi(\overline{x},y)$. Let $(p^{\psi},\widehat{\Lambda}^{\psi},\widehat{\psi},\widehat{L}^{\psi}(u))$, $(\Lambda^{\xi},\xi,L^{\xi})$ as it was done in Case 1 in the proof of Lemma 10. Again, instead of computing these values, each node receives them as certificates, and checks them. The values of $p,p^{\psi},\widehat{\Lambda}^{\psi}$, $\widehat{\psi}$, $\Lambda^{\xi}$, and $\xi$ do not depend on the input graph, and can be computed by each node. Then, each node can construct the input values (1) to (4) given in Definition 4 with respect to to $\xi$ and $\psi$. By the induction hypothesis, the correctness of $\widehat{L}^{\psi}(u)$ at each node $u$ can be certified with $\mathcal{O}(\log n)$-bit certificates.
Recall that, in Lemma 10, the value of $L^{\xi}(u)$ on each node $u$ was computed using Lemma 9, without any communication, under the assumption that each node knows a encoding a $p^{\psi}$-skeleton $\tau_{p^{\psi}}$ of $G$. The prover provides the encoding of a $p^{\psi}$-skeleton to each node as a $\mathcal{O}(\log n)$-bit certificate, which can be checked using Lemma 17. Finally, the correctness of $\widehat{L}$ can be certified with $\mathcal{O}(\log n)$-bit certificates, using the scheme of Lemma 18 with given values $(\Lambda^{\xi},(\exists y,\xi(\overline{x},y)),L^{\xi})$ and $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{L})$ and $\tau_{p}$. ∎
5.4 Proof of Theorem 4
We are now ready to prove the main result of this section.
See 4
@@ 5.4 Proof of Theorem 4
Let $\varphi\in\text{FO}[\Lambda]$. We proceed the same way as in Theorem 1. Specifically, let us redefine $\varphi(x)$, as $\varphi$ with a dummy free-variable $x$. The prover provides each node with the following certificate:
@@
    a positive integer $p$, a set $\widehat{\Lambda}$, a $p$-lca-reduced formula $\widehat{\varphi}$, and a label set $\widehat{L}(u)$;
@@
    the certificate of $u$ used to certify the $p$-skeleton $\tau_{p}$ of $G$;
@@
    the certificate $\textsf{Cert}(u)$ for certifying that $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{L})$ is the reduction of $(\Lambda,\varphi,\tau_{p},L)$.
The part (i) of the certificate can be encoded in $\mathcal{O}(1)$ bits, the part (ii) can be encoded in $\mathcal{O}(\log n)$ bits, and Lemma 19 asserts that a certificate for (iii) exists and, furthermore, it can be encoded in $\mathcal{O}(\log n)$ bits.
The verifier proceeds as follows. Each node checks that $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{L})$ is the reduction of $(\Lambda,\varphi,\tau_{p},L)$ using the certificates given in (iii) using the verifier $\mathcal{A}$ in Lemma 19, with certificates (i) and (ii) provided to the nodes. If a node rejects in $\mathcal{A}$, then the node rejects, otherwise it carries on the verification as described hereafter.
Assuming that no nodes have rejected so far, we have that $\textsf{true}(\varphi,G,L)=\textsf{true}(\widehat{\varphi},G,\widehat{L})$. Since $\varphi$ has only one dummy variable, the truth value of $\varphi(x)$, as well as the one of $\widehat{\varphi}(x)$, is the same for every node in $G$. Moreover, the formula $\widehat{\varphi}(x)$ is a Boolean combination of label predicates of $x$. Therefore, every node $u$ can check $\widehat{\varphi}(u)$ using $\widehat{L}(u)$, and it accepts if $\widehat{\varphi}(u)$ holds, and rejects otherwise.
@@
For establishing completeness of the certification scheme, let us suppose that $G\models\varphi$. Then, by Lemma 10, we have that $\varphi$ is reducible in $\mathcal{G}$. Then there exist certificates for (i) and (ii), and, by Lemma 19, there exist certificates for (iii) such that every node accepts when running verifier $\mathcal{A}$. Since $G\models\varphi$ we have that $\textsf{true}(\varphi,G)=\textsf{true}(\widehat{\varphi},G,\widehat{L})=V(G)$. It follows that $\widehat{\varphi}(u)$ is true at every node $u$, and thus every node accepts.
For establishing soundness, let us suppose that $G\not\models\varphi$, and let us suppose that the nodes received certificates (i)-(iii) such that every node accepts when renning the verifier $\mathcal{A}$. Then $\textsf{true}(\varphi,G,L)=\textsf{true}(\widehat{\varphi},G,\widehat{L})$. However, in this case $\textsf{true}(\varphi,G,L)=\emptyset$. Therefore $\widehat{\varphi}(u)$ does not hold at any node $u$, and thus all nodes reject. ∎
@@ 6 Lower Bounds and Impossibility Results
In this section we give evidence that our results somehow represent the limit of tractabilty of distributed model checking and distributed certification of properties on graphs of bounded expansion.
@@ 6.1 No Extension to Monadic Second-Order Logic
The proof of [46] is consists in a reduction to $\textsf{Disj}_{n^{2}}$. In the reduction each instance $(A,B)$ of $\textsf{Disj}_{n^{2}}$ is associated with a $O(n)$-node graph $G_{A,B}$. The set of vertices of $G_{A,B}$ is divided in two sets of vertices, denoted $V_{A}$ and $V_{B}$, and satisfies the following conditions:
@@
The construction is also useful for a lowerbound in the CONGEST model. Let $\mathcal{A}$ be an algorithm deciding the membership in Non-3-Coloring in $\mathcal{R}(n)$ rounds. Then Alice and Bob construct $G_{A,B}$ as above, and simulate the $\mathcal{R}(n)$ rounds of $\mathcal{A}$ on it, communicating on each round to other player the messages transmitted over the edges $E_{A,B}$. The players accept if all the nodes of their sides accept. On each communication round the players send $\mathcal{O}(\log n)$ bits per edge of $E_{A,B}$. Hence, the nodes interchange $\mathcal{O}(R(n)\cdot|E_{A,B}|\cdot\log n)=\mathcal{O}(R(n)\cdot\log^{2}n)$ bits in total. From Proposition 6 we obtain that $\mathcal{O}(R(n)\cdot\log^{2}n)=D(\textsf{Disj}_{n^{2}})=\Omega(n^{2})$. Therefore, $R(n)=\Omega(n^{2}/\log^{2}n)$.
@@
Theorem 5 implies that the difficulty of certifying MSO properties over graphs of max degree 5 is (up to logarithmic factors) as hard as the hardest problem over that family. Indeed, a graph of max degree $5$ can be encoded using $\mathcal{O}(n\log n)$ bits, e.g. by adjacency lists of all nodes. We can use such encoding to design a certification scheme that certifies any property on graphs of maximum degree $5$: every node interprets its certificate as an encoding of a graph of maximum degree $5$, and checks that every neighbor received the same encoding, and that its local view corresponds with the one in the encoding. If the conditions are verified, the nodes can compute any property of the input graph without any further communication.
@@
Similar to the proof of Theorem 5 we use the construction of Göös and Suomella and the reduction of Garey, Johnson and Stockmeyer for planar graphs. Given an $n$-node graph $G$, the reduction of [38] for planar graphs first pick an arbitrary embedding of $G$ in the plane. In this embedding, there might be a number $\textsf{cr}(G)$ of edge crossings (they must exist if $G$ is non-planar). Then, a planar graph $\widetilde{G}$ is constructed from $G$ by replacing on point where a pair of edges cross, by the crossing gadget shown in Figure 3. This gadget allows to transmit the coloring of one of the endpoints of the edge in a planar way. For each crossing, we add 13 new vertices. In the worst case, $\textsf{cr}(G)=\Theta(n^{4})$. Hence, the obtained graph has $\widetilde{G}$ has $\Theta(n^{4})$ nodes.
@@
Let $\mathcal{G}$ be the class of planar graphs. As we did for Theorem 5, we plug the reduction construction of Garey, Johnson and Stockmeyer to the construction of Göös and Suomella. More precisely, given an instance $(A,B)$ of $\textsf{Disj}_{n}$, we build graph $G_{A,B}$ (c.f. proof of Theorem 5). Then, using the crossing gadget we obtain a planar graph $\widetilde{G}_{A,B}$. We partition the nodes of $\widetilde{G}_{A,B}$ into $\widetilde{V}_{A}$ and $\widetilde{V}_{B}$. All the nodes inside gadgets created for edges with both endpoint in $V_{A}$ (respectively $V_{B}$) belong to $\widetilde{V}_{A}$ (respectively $\widetilde{V}_{B}$). We also consider that the nodes inside the gadgets used for the edges in the set $E_{A,B}$ of $G_{A,B}$ belong to $V_{A}$. We obtain that $\widetilde{G}_{A,B}$ satisfies conditions (C1)-(C4).
The remaining of the proof is analogous to Theorem 5. Let $\mathcal{P}$ be a certification scheme for Non-3-Coloring with certificates of size $c(n)$ under the promise that the input graph is planar. On input $A,B\subseteq[n]$, Alice and Bob construct $\widetilde{G}_{A,B}$ and simulate $\mathcal{P}$ in the same way that we explained for $G_{A,B}$. Since $\widetilde{G}_{A,B}$ is a graph with $\Theta(n^{4})$ nodes, we obtain that $\mathcal{O}(c(n^{4})\cdot\log n)=N(\textsf{Disj}_{n^{2}})=\Omega(n^{2})$, hence $c(n^{4})=\Omega(n^{2}/\log n)$, and then $c(n)=\Omega(\sqrt{n}/\log n)$. ∎

### v2

@@ 6.3 Certifying Reductions on Graphs of Bounded Expansion
Definition 6.6.
Let $\varphi\in\mathrm{FO}[\Lambda]$. For every labeled graph $(G,\ell)$ with $G\in\mathcal{G}$, let us assume that every node $u$ of $G$ is provided with the following values:
@@
    The set $\Lambda$, the formula $\varphi$, and the label $\ell(u)\subseteq\Lambda$;
@@
    A set $\widehat{\Lambda}$, a formula $\widehat{\varphi}$, and a label ^$\widehat{\ell}(u)\subseteq\widehat{\Lambda}$;
@@
    The encoding of a $p$-skeleton $S$ of $G$.
We say that $\varphi$ admits a distributedly certifiable reduction on $\mathcal{G}$ if there is a certification scheme using $\mathcal{O}(\log n)$-bits certificates for certifying that $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ is the reduction of $(\Lambda,\varphi,S,\ell)$ on $G$.
In a way similar to what was done in Section 4, we show that every first-order formula over labeled graphs admits a certifiable reduction. We fist establish this result for existential formulas.
Lemma 6.7.
Every existential formula in $\mathrm{FO}[\Lambda]$ admits a distributed certifiable reduction on $\mathcal{G}$.
@@
Let $\varphi\in\mathrm{FO}[\Lambda]$ be an existential formula. Consider that each node is provided with the values listed in Definition 6.6. Let $\varphi(\overline{x})=\exists\overline{y}\,\psi(\overline{x},\overline{y})$, where $\psi$ is quantifier free. Every node checks that $p=|\overline{x}|+|\overline{y}|$. The scheme of Lemma 6.5 enables to certify that the encoding $S$ correctly defines a $p$-skeleton of $G$. Let $C_{p}$ be the $(p,f(p))$-treedepth coloring that defines $S$. The same construction as in Lemma 4.3 can then be used. More precisely, for every $U\in\textsf{Col}(p)$, let $\widetilde{\Lambda}^{U}$, $\widetilde{\psi}^{U}$, ~$\widetilde{\ell}^{U}$, $\Lambda^{U}$, $\psi^{U}$, and $\ell^{U}$ be defined as in the proof of Lemma 4.3. Instead of computing these values, each node receives them from the prover as certificates. The values of $\widehat{\Lambda}$, $\widehat{\varphi}$, $\widetilde{\Lambda}^{U}$, $\widetilde{\psi}^{U}$, $\Lambda^{U}$, and $\psi^{U}$ do not depend on the input graph, and can be computed by each node. The labeling ~$\tilde{\ell}^{U}$ can be computed from $\ell$ by a 1-round algorithm in the BROADCAST CONGEST model. Hence, using Lemma 6.2, the correctness of each labeling ~$\tilde{\ell}^{U}$ can be certified with $\mathcal{O}(\log n)$-bit certificates. The correctness of each labeling ^$\widehat{\ell}^{U}$ can be certified using Lemma 6.3, using $\mathcal{O}(\log n)$-bit certificates. Finally, each node can check the correctness of $\widehat{\ell}$, and accept or reject accordingly. ∎
@@
Lemma 6.8.
Every formula in $\mathrm{FO}[\Lambda]$ admits a distributed certifiable reduction on $\mathcal{G}$.
@@
Let $\varphi\in\mathrm{FO}[\Lambda]$ be an existential formula. Let us assume that each node is provided with the values described in Definition 6.6, i.e. the tuples $(\Lambda,\varphi,\ell)$ and $(p,\hat{\Lambda},\hat{\varphi},\hat{\ell})$, and the encoding of a $p$-skeleton $S$. We follow the same induction as in Lemma 4.3, over the quantifier depth of $\varphi$. The base case is trivial as the construction requires no communication. Furthermore, if $\varphi=\neg\psi$ where $\psi$ admits a distributed certifiable reduction, then $\varphi$ admits a distributed certifiable reduction as well, as we described in Lemma 4.3.
It remains to consieder the case $\varphi(\overline{x})=\exists y\,\psi(\overline{x},y)$. Let $(p^{\psi},\widehat{\Lambda}^{\psi},\widehat{\psi},\widehat{\ell}^{\psi}(u))$, $(\Lambda^{\xi},\xi,\ell^{\xi})$ as it was done in Case 1 in the proof of Lemma 4.5. Again, instead of computing these values, each node receives them as certificates, and checks them. The values of $p,p^{\psi},\widehat{\Lambda}^{\psi}$, $\widehat{\psi}$, $\Lambda^{\xi}$, and $\xi$ do not depend on the input graph, and can be computed by each node. Then, each node can construct the input values (1) to (4) given in Definition 6.6 with respect to to $\xi$ and $\psi$. By the induction hypothesis, the correctness of ^$\widehat{\ell}^{\psi}(u)$ at each node $u$ can be certified with $\mathcal{O}(\log n)$-bit certificates.
Recall that, in Lemma 4.5, the value of $\ell^{\xi}(u)$ on each node $u$ was computed using Lemma 4.4, without any communication, under the assumption that each node knows a encoding a $p^{\psi}$-skeleton $S$ of $G$. The prover provides the encoding of a $p^{\psi}$-skeleton to each node as a $\mathcal{O}(\log n)$-bit certificate, which can be checked using Lemma 6.5. Finally, the correctness of $\widehat{\ell}$ can be certified with $\mathcal{O}(\log n)$-bit certificates, using the scheme of Lemma 6.7 with given values $(\Lambda^{\xi},(\exists y,\xi(\overline{x},y)),\ell^{\xi})$ and $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ and $S$. ∎
6.4 Proof of Theorem 1.5
We are now ready to prove Theorem 1.5.
@@
Let $\varphi\in\mathrm{FO}[\Lambda]$. We proceed the same way as in Theorem 1.2. Specifically, let us redefine $\varphi(x)$, as $\varphi$ with a dummy free-variable $x$. The prover provides each node with the following certificate:
@@
    a positive integer $p$, a set $\widehat{\Lambda}$, a $p$-lca-reduced formula $\widehat{\varphi}$, and a label set ^$\widehat{\ell}(u)$;
@@
    the certificate of $u$ used to certify the $p$-skeleton $S$ of $G$;
@@
    the certificate $\textsf{Cert}(u)$ for certifying that $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ is the reduction of $(\Lambda,\varphi,S,\ell)$.
The part (i) of the certificate can be encoded in $\mathcal{O}(1)$ bits, the part (ii) can be encoded in $\mathcal{O}(\log n)$ bits, and Lemma 6.8 asserts that a certificate for (iii) exists and, furthermore, it can be encoded in $\mathcal{O}(\log n)$ bits.
The verifier proceeds as follows. Each node checks that $(p,\widehat{\Lambda},\widehat{\varphi},\widehat{\ell})$ is the reduction of $(\Lambda,\varphi,S,\ell)$ using the certificates given in (iii) using the verifier $\mathcal{A}$ in Lemma 6.8, with certificates (i) and (ii) provided to the nodes. If a node rejects in $\mathcal{A}$, then the node rejects, otherwise it carries on the verification as described hereafter.
Assuming that no nodes have rejected so far, we have that $\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},G,\widehat{\ell})$. Since $\varphi$ has only one dummy variable, the truth value of $\varphi(x)$, as well as the one of $\widehat{\varphi}(x)$, is the same for every node in $G$. Moreover, the formula $\widehat{\varphi}(x)$ is a Boolean combination of label predicates of $x$. Therefore, every node $u$ can check $\widehat{\varphi}(u)$ using ^$\widehat{\ell}(u)$, and it accepts if $\widehat{\varphi}(u)$ holds, and rejects otherwise.
@@
For establishing completeness of the certification scheme, let us suppose that $G\models\varphi$. Then, by Lemma 4.5, we have that $\varphi$ is reducible in $\mathcal{G}$. Then there exist certificates for (i) and (ii), and, by Lemma 6.8, there exist certificates for (iii) such that every node accepts when running verifier $\mathcal{A}$. Since $G\models\varphi$ we have that $\textsf{true}(\varphi,G)=\textsf{true}(\widehat{\varphi},G,\widehat{\ell})=V(G)$. It follows that $\widehat{\varphi}(u)$ is true at every node $u$, and thus every node accepts.
For establishing soundness, let us suppose that $G\not\models\varphi$, and let us suppose that the nodes received certificates (i)-(iii) such that every node accepts when renning the verifier $\mathcal{A}$. Then $\textsf{true}(\varphi,G,\ell)=\textsf{true}(\widehat{\varphi},G,\widehat{\ell})$. However, in this case $\textsf{true}(\varphi,G,\ell)=\emptyset$. Therefore $\widehat{\varphi}(u)$ does not hold at any node $u$, and thus all nodes reject. ∎
@@ 7 Lower Bounds and Impossibility Results
In this section we give evidence that our results somehow represent the limit of tractability of distributed model checking and distributed certification of properties on graphs of bounded expansion.
@@ 7.1 No Extension to Monadic Second-Order Logic
The proof of [36] is consists in a reduction to $\textsf{Disj}_{n^{2}}$. In the reduction each instance $(A,B)$ of $\textsf{Disj}_{n^{2}}$ is associated with a $\mathcal{O}(n)$-node graph $G_{A,B}$. The set of vertices of $G_{A,B}$ is divided in two sets of vertices, denoted $V_{A}$ and $V_{B}$, and satisfies the following conditions:
@@
The construction is also useful for a lowerbound in the CONGEST model. Let $\mathcal{A}$ be an algorithm deciding the membership in Non-3-Coloring in $\mathcal{R}(n)$ rounds. Then Alice and Bob construct $G_{A,B}$ as above, and simulate the $\mathcal{R}(n)$ rounds of $\mathcal{A}$ on it, communicating on each round to other player the messages transmitted over the edges $E_{A,B}$. The players accept if all the nodes of their sides accept. On each communication round the players send $\mathcal{O}(\log n)$ bits per edge of $E_{A,B}$. Hence, the nodes interchange $\mathcal{O}(R(n)\cdot|E_{A,B}|\cdot\log n)=\mathcal{O}(R(n)\cdot\log^{2}n)$ bits in total. From Proposition 7.1 we obtain that $\mathcal{O}(R(n)\cdot\log^{2}n)=D(\textsf{Disj}_{n^{2}})=\Omega(n^{2})$. Therefore, $R(n)=\Omega(n^{2}/\log^{2}n)$.
@@
Theorem 7.2 implies that the difficulty of certifying MSO properties over graphs of max degree 5 is (up to logarithmic factors) as hard as the hardest problem over that family. Indeed, a graph of max degree $5$ can be encoded using $\mathcal{O}(n\log n)$ bits, e.g. by adjacency lists of all nodes. We can use such encoding to design a certification scheme that certifies any property on graphs of maximum degree $5$: every node interprets its certificate as an encoding of a graph of maximum degree $5$, and checks that every neighbor received the same encoding, and that its local view corresponds with the one in the encoding. If the conditions are verified, the nodes can compute any property of the input graph without any further communication.
@@
Similar to the proof of Theorem 7.2 we use the construction of Göös and Suomella and the reduction of Garey, Johnson and Stockmeyer for planar graphs. Given an $n$-node graph $G$, the reduction of [30] for planar graphs first pick an arbitrary embedding of $G$ in the plane. In this embedding, there might be a number $\textsf{cr}(G)$ of edge crossings (they must exist if $G$ is non-planar). Then, a planar graph $\widetilde{G}$ is constructed from $G$ by replacing on point where a pair of edges cross, by the crossing gadget shown in Figure 4. This gadget allows to transmit the coloring of one of the endpoints of the edge in a planar way. For each crossing, we add 13 new vertices. In the worst case, $\textsf{cr}(G)=\Theta(n^{4})$. Hence, the obtained graph has $\widetilde{G}$ has $\Theta(n^{4})$ nodes.
@@
Let $\mathcal{G}$ be the class of planar graphs. As we did for Theorem 7.2, we plug the reduction construction of Garey, Johnson and Stockmeyer to the construction of Göös and Suomella. More precisely, given an instance $(A,B)$ of $\textsf{Disj}_{n}$, we build graph $G_{A,B}$ (c.f. proof of Theorem 7.2). Then, using the crossing gadget we obtain a planar graph $\widetilde{G}_{A,B}$. We partition the nodes of $\widetilde{G}_{A,B}$ into $\widetilde{V}_{A}$ and $\widetilde{V}_{B}$. All the nodes inside gadgets created for edges with both endpoint in $V_{A}$ (respectively $V_{B}$) belong to $\widetilde{V}_{A}$ (respectively $\widetilde{V}_{B}$). We also consider that the nodes inside the gadgets used for the edges in the set $E_{A,B}$ of $G_{A,B}$ belong to $V_{A}$. We obtain that $\widetilde{G}_{A,B}$ satisfies conditions (C1)-(C4).
The remaining of the proof is analogous to Theorem 7.2. Let $\mathcal{P}$ be a certification scheme for Non-3-Coloring with certificates of size $c(n)$ under the promise that the input graph is planar. On input $A,B\subseteq[n]$, Alice and Bob construct $\widetilde{G}_{A,B}$ and simulate $\mathcal{P}$ in the same way that we explained for $G_{A,B}$. Since $\widetilde{G}_{A,B}$ is a graph with $\Theta(n^{4})$ nodes, we obtain that $\mathcal{O}(c(n^{4})\cdot\log n)=N(\textsf{Disj}_{n^{2}})=\Omega(n^{2})$, hence $c(n^{4})=\Omega(n^{2}/\log n)$, and then $c(n)=\Omega(\sqrt{n}/\log n)$. ∎